Circular No. 5613

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 C. Telesco, Marshall Space Flight Center (MSFC), NASA; R. Pina,

University of California at San Diego; S. Fajardo, State University of New York; G. Fishman and C. Kouveliotou, MSFC; and J. van Paradijs, University of Amsterdam, report: “We have detected the γ-ray transient GRO J0422+32 (cf. IAUC 5580) at 10.8 μm using the MSFC 20-pixel bolometer array at the 3-m Infrared Telescope Facility on Mauna Kea. The 10.8-μm flux density was 51 ± 9 mJy (detected S/N = 5.7; the uncertainty does not include ± 7 percent in absolute calibration). Only the pixel centered on the optical counterpart (IAUC 5588) showed statistically significant S/N. Total on-source integration time was 1.17 hr, distributed uniformly throughout the interval Sept. 9.52-9.58 UT.”

P. Garnavich, Dominion Astrophysical Observatory; and A. Noriega-Crespo, University of Washington, communicate: “We obtained CCD photometry of GRO J0422+32 on five nights between Aug. 17 and Sept. 7, using the 0.8-m telescope at the Manastash Ridge Observatory of the University of Washington. On four of the nights, images were taken using a B filter and 30- to 40-s integrations. The power spectra of the light curves are flat from 1 to ∼ 15 mHz. Irregular variations of 5-10 percent in brightness are seen at frequencies lower than ∼ 1 mHz. On Sept. 2, a series of unfiltered CCD images were obtained with 3.0-s integrations and 8.4 s between mid-exposures. The light curve spans 1.5 hr. The power spectrum shows no evidence of a peak near the frequency of the 35-mHz QPO seen at x-ray wavelengths (IAUC 5608). We find, however, that the variance in the brightness estimated from individual exposures is 5 times larger than that expected from the combination of CCD read-noise and photon shot-noise. Photometry of other stars on the CCD images confirms the surprisingly large variance. We conclude that the power spectrum of the GRO J0422+32 light curve is ‘white’ from 1 to 50 mHz and that the amplitude measured as a standard deviation about a mean brightness is ∼ 0.05 magnitude in unfiltered light. A closer examination of our low-time-resolution, B-filter data also shows excess noise, but with a standard deviation of < 0.03 magnitude.”

Further visual magnitude estimates by P. Schmeer, Bischmisheim, Germany (cf. IAUC 5597): Aug. 29.08 UT, 12.9; Sept. 7.13, 12.7.

1992 September 17 Daniel W. E. Green